OpenStudy (anonymous):

Evaluate the function g(t) = 3/2t - 5/2, when t = 5

OpenStudy (cggurumanjunath):

put t =5 in g(t)

OpenStudy (cggurumanjunath):

find g(5)=g(t)

OpenStudy (anonymous):

\[g(t) =\frac {3}{2t} - \frac{5}{2}\] \[g(5) =\frac {3}{25} - \frac{5}{2}\] Find a common denominator and solve.

OpenStudy (cggurumanjunath):